OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x - y - 2z if x = -2, y = 3, and z = -2

OpenStudy (anonymous):

-15 15 -7 7

OpenStudy (anonymous):

plug everything in and its simple. 4 x -2= -8 -8-(3)-2(-2) -8-3+4 -11+4=-7 answer= -7

OpenStudy (texaschic101):

I agree :)